If x^2 = 1-x, find the value of the integers a and b if x^4 = ax + b
Answer by rothauserc(4718)

x^4 = ax +b
:
x^2 * x^2 = ax +b
:
(1-x) * (1-x) = ax +b
:
1 -2x +x^2 = ax +b
:
1 -2x +1 -x = ax +b
:
-3x +2 = ax +b
:
***********************
a = -3 and b = 2
***********************
: